OpenStudy (anonymous):

The length of a rectangular boat storage shed must be 4 meters more than width and the perimeter must be st least 100 meters. what is the range of values for the witdth 23 meters exactly at least 23 meters at most 23 meters ???

OpenStudy (mathmale):

I'd suggest you draw a diagram. Label the length and width with " l " and " w " respectively. If the length must be 4 meters more than the width, how would you represent that algebraically? Please write a formula for " l " in terms of w. What's the formula for the perimeter of a rectangle of length " l " and width " w "?

OpenStudy (anonymous):

Im a returning student after 10 plus years so everything is really rusty give me a sec

OpenStudy (anonymous):

4+w=x

OpenStudy (mathmale):

Looks good! Remember how to find the perimeter of a rectangle?

OpenStudy (anonymous):

unfortunately not

OpenStudy (mathmale):

do you have a texbook, or are you using the web as your primary source of information in this problem solving?

OpenStudy (anonymous):

the web

OpenStudy (mathmale):



OpenStudy (mathmale):

I'd suggest you Google terms such as "perimeter." That'd give you definitions in a great hurry, although you'd still have to wade through a lot of web pages to get the definition you need. Does my diagram give you any idea of how to find the perimeter? We have two lengths, each measured by w+4, and two widths, each measured by w, and we have to add up all these lengths to obtain and expression for perimeter. P=2w + 2x, where w=width and x=length.

OpenStudy (anonymous):

it helps

OpenStudy (mathmale):

Our math problem states that the perimeter has to be at least 100 m, or greater. We have to translate that into an algebraic inequality: P => 100 m. that " => " is symbolic for "greater than or equal to."

OpenStudy (mathmale):

P=2w + 2x, where w=width and x=length. Write the appropriate inequality.

OpenStudy (anonymous):

why this one gotta be so difficult

OpenStudy (mathmale):

It'll become easier as you get back into the swing of things academically. If P=>100 m., then 2w+2x => 100 m.

OpenStudy (anonymous):

crap thats what i was writting lol

OpenStudy (mathmale):

If 2w + 2x => 100 m, and x=4+w (which we have from earlier), then 2w + 2(4+w) => 100 m. All I did was to substitute 4+w for x.

OpenStudy (mathmale):

What do you think the next step of the problem solution is?

OpenStudy (anonymous):

2w+8+2w=100m -- 4w+8=100m right track ??

OpenStudy (anonymous):

w=23 ?!?!?!?

OpenStudy (anonymous):

so the answer is exactly 23 meters ;-)

OpenStudy (anonymous):

correct ??

OpenStudy (mathmale):

As far as you go, yes! But remember, we're dealing with an inequality here. So, where you have w=23 m., you should actually have w => 23 m. Please translate that into words.

OpenStudy (anonymous):

the width is greater than

OpenStudy (mathmale):

the width is greater than or equal to 23 meters.

OpenStudy (anonymous):

the equal sign duh

OpenStudy (mathmale):

I'll be happy to help you with other problems. Basically, you need to become familiar again with the language of algebra. When in doubt, please look up the vocabulary. You could do that by simply Googling the term in question, or by going to dictionary.com.

OpenStudy (mathmale):

Again: => represents "greater than or equal to."

OpenStudy (anonymous):

so the answer is at most i would say

OpenStudy (anonymous):

23 meters

OpenStudy (mathmale):

The length of a rectangular boat storage shed must be 4 meters more than width and the perimeter must be st least 100 meters. what is the range of values for the witdth 23 meters? exactly 23 meters at least 23 meters at most 23 meters think again: How does w => 23 m translate into English?

OpenStudy (anonymous):

least

OpenStudy (mathmale):

Need to translate into English EVERYTHING you see in w => 23 m: "w is equal to or greater than 23 meters." Alternatively, "w is at least 23 meters." You were on the right track, but you typed in only one word: least; you need all of "w is at least 23 meters."

OpenStudy (anonymous):

may I ask you another question

OpenStudy (mathmale):

Yes. I'll need to stop after that, but yes. We can certainly meet later on as well.

OpenStudy (mathmale):

Please post the new question in the usual way, not here.